OpenStudy (anonymous):

why does the reciprocal of (2√5)/5 equals √5/2 ???

hartnn (hartnn):

because \(\huge \frac{\sqrt 5}{5}=\frac{\sqrt 5}{\sqrt5.\sqrt 5}=\frac{1}{\sqrt 5}\)

OpenStudy (anonymous):

i still dont get it..><

hartnn (hartnn):

\(\huge \frac{2\sqrt 5}{5}=\frac{2\sqrt 5}{\sqrt5.\sqrt 5}=\frac{2}{\sqrt 5}\) which step ?

OpenStudy (anonymous):

(2√5)/5 can also be written as \[\frac{ 2 }{ \sqrt{5} }\]

hartnn (hartnn):

reciprocal means inversing the ratio, a/b becomes b/a

OpenStudy (anonymous):

why is the denominator √5 multiply by √5

hartnn (hartnn):

any number can be expressed as \(a=\sqrt a.\sqrt a\)

hartnn (hartnn):

\(x^m.x^n=x^{m+n} \\so, \huge \sqrt 5.\sqrt 5= 5^{1/2}.5^{1/2}=5^{1/2+1/2}=5^1\)

hartnn (hartnn):

doubts ?

OpenStudy (anonymous):

wait basically....√5 and 5 is the same?

hartnn (hartnn):

no.... 5 and \(\sqrt 5.\sqrt 5\) are same.

OpenStudy (anonymous):

but why do we have to multply by √5??

OpenStudy (anonymous):

OMG i think i got it!!

hartnn (hartnn):

i am glad!

OpenStudy (anonymous):

so first, the reciprocal of (2√5)/5 is 5/2√5

OpenStudy (anonymous):

but 5 = √5 . √5

hartnn (hartnn):

es, you can do that way also!

OpenStudy (anonymous):

and √5 crosses out which equals. √5 / 2!!

OpenStudy (anonymous):

thank you soooo much!!!! im sorry if i put some streess on you!! i am really bad at math ><

hartnn (hartnn):

no problem at all :) welcome ^_^